# Predicting the single-site and multi-site event discrimination power of dual-phase time projection chambers

### v2

@@ Abstract
Dual-phase xenon time projection chambers (TPCs) are widely used in searches for rare dark matter and neutrino interactions, in part because of their excellent position reconstruction capability in 3D. Despite their millimeter-scale resolution along the charge drift axis, xenon TPCs face challenges in resolving single-site (SS) and multi-site (MS) interactions in the transverse plane. In this paper, we build a generic TPC model with an idealized light-based signal readout, and use Fisher Information (FI) to study its theoretical capability of differentiating SS and MS events. We also demonstrate via simulation that this limit can be approached with conventional reconstruction algorithms like maximum likelihood estimation, and with a convolutional neural network classifier. The implications of this study on future TPC experiments will be discussed.
@@ 1 Introduction
The time projection chamber (TPC) [1] is one of the most sensitive particle detection technologies. A TPC is a drift chamber that transports ionization electrons produced by particle interaction across the detector volume with an electric field; at one end of the detector the electrons are collected with a position sensitive charge sensing device. The charge sensor provides two-dimensional (2D) position information in the collection plane perpendicular to the electric field. In addition, if the interaction time is known, for example, via an accelerator beam signal or prompt scintillation accompanying ionization, the third coordinate can be reconstructed from the electron drift time, enabling full three-dimensional (3D) event reconstruction. For gas TPCs that produce relatively long particle tracks, the track ionization density can be measured, allowing the particle type to be determined. For scintillating TPCs, the charge-to-scillation ratio may also help identify different particle interactions [2]. In addition, combining the scintillation and ionization energy channels enables an accurate energy reconstruction for the interaction by reducing the inter-channel fluctuation [3].
TPCs using liquid argon and xenon, which combine the benefits of a bright scintillator, a large target mass, and low electron affinity, are widely used in the search for rare particle interactions such as those expected from neutrinos and dark matter [4, 5, 6, 7, 8, 9]. A dual-phase noble liquid TPC further enhances the TPC’s sensitivity to low-energy interactions by introducing a gas layer above a liquid target. When the electrons are drifted into the gas layer, they can produce a large number of electroluminescence photons under a strong electric field. Therefore, both the scintillation and the ionization signals can be conveniently collected with a single set of photosensors. Typically, a single extracted electron can produce hundreds of electroluminescence photons, allowing signals as small as single ionization electrons to be detected with high efficiency [10]. Dual-phase noble-liquid TPCs, especially those with a xenon target, have been widely used in dark matter search experiments that primarily focus on very low-energy signals. Currently, there are three xenon TPCs with multi-ton target masses are being operated, including the LZ experiment [4] at the Sanford Underground Research Facility (SURF), the XENONnT experiment [11] at Gran Sasso National Laboratory (LNGS), Italy, and the PandaX-nT experiment [8] at the Jinping Underground Laboratory in China. These detectors have been leading the searches for a range of hypothetical dark matter interactions, and their successes have inspired even larger detectors such as XLZD [12] and PandaX-20T to be proposed.
@@
One important application of SS/MS discrimination is in the search for neutrino-less double beta decays ($0\nu\beta\beta$) of the isotope 136Xe in a dual-phase xenon TPC [15]. $0\nu\beta\beta$ is predicted to occur in some double beta decay isotopes if neutrinos are Majorana fermions. A 136Xe $0\nu\beta\beta$ signal consists of two $\beta$ electrons whose energies sum to exactly $Q_{\beta\beta}=2.457$ MeV, while the primary background is induced by $\gamma$ rays from trace amounts of natural radioactivity in the detector materials. In liquid xenon, the $\beta$ electrons are stopped within millimeters and the events appear point-like, while the gamma rays tend to Compton scatter and deposit energy in multiple locations separated by several centimeters. Reconstruction of event topology therefore provides a powerful means of background discrimination. The nEXO experiment proposes to search for $0\nu\beta\beta$ in 136Xe with $\sim$5 tons of enriched xenon in a single-phase TPC and is expected to fully cover the inverted hierarchy region of neutrino mass with 10 years of operation. Because of its large total mass (60–80 ton), the dual-phase XLZD experiment would contain a comparable mass of 136Xe to that of nEXO and its sub-percent energy resolution near the 136Xe $Q$-value enables it to carry out a sensitive $0\nu\beta\beta$ search [16]. In both cases, 214Bi decay gamma rays are expected to be the dominant background in $0\nu\beta\beta$ energy region even with meticulous material cleanliness protocols [15, 16]. Therefore, a thorough understanding of the dual-phase xenon TPC’s capability in separating SS and MS events is critical.
@@
2 Detector model and sensitivity predictions
We develop a generic dual-phase TPC model with a light-based readout for ionization signals to study its theoretical performance in reconstructing SS and MS events across different event topologies as a function of the photosensor size and the total detected photon count. Photosensor diameters of 3", 2", 1", 15 mm, and 6 mm are considered, which represents the sizes of commonly used PMTs and SiPMs in dual-phase TPCs.
@@ 2.1 Detector model
Figure 1: Left - an illustration of the generic TPC model used in our studies; Right - a cross-section view of the TPC model detailing the photosensor array configuration; the asterisks represent two light emission sites separated by a distance $d$ and $r_{1}$ and $r_{2}$ represent the radial distances between the light sources and the PMT under investigation.
@@
The detector model, as illustrated in Fig. 1, is simplified to capture the essential features of a generic dual-phase TPC. The electroluminescence region is modeled as a 1 cm xenon gas layer above the liquid–gas interface, where the electric field is assumed to be perfectly perpendicular to the liquid surface. With this approximation, electroluminescence photons excited by extracted electrons at a certain xyx-y location are produced uniformly along the vertical drift paths in this gap. Our model omits physical TPC electrodes, which generate the TPC drift field but may introduce field non-uniformities near the grid wires and cause shadowing effect in the light distribution. An array of photonsensors is placed 2 cm above the liquid surface to create a 1 cm separation between the upper end of photon tracks and the nearest sensor. This light source-sensor separation is smaller than those used in large TPCs, but is chosen to allow the studies of small sensors with the same TPC configuration; the effects of the sensor separation is discussed in Sec. 4.2 and a study using the LUX detector geometry can be found in Appendix B. Each sensor is assumed to be circular in shape  [^1] SiPMs are often fabricated in square shapes, which can be modeled using the same method as described in this section; we model all sensors in the same shape to directly compare their performances. in a hexagonal close-packed configuration, yielding a photocathode fill factor of $\sim$91%. The sensor array is assumed to be sufficiently large to capture essentially all photons that could be collected, simulating the scenario of near-center events in a large TPC. While partial photon collection at detector edges could be implemented via detailed geometry, such effects are not considered in this theoretical study. In practice, Fresnel reflections at large incident angles, combined with the small vertical separation between light source and sensor, limits the acceptance of photons from large radial distances. We determine that a 40 cm array is adequate for studies where photons are emitted from the center. Studies requiring light production far away from the center use larger detector and sensor sizes accordingly.
@@
A Geant4-based Monte Carlo simulation of the optical process is used to predict the photon detection pattern in the photosensor array for electroluminescence light emitted from a specific xyx-y location. In the simulation, all material surfaces, except the liquid-gas boundary, are assumed to be photon-absorbing so that no reflectance of light is in effect except at the liquid surface. Despite the simplification, our simulation approximately reproduces the radial photon distribution in the LUX experiment [14] when the same sensor placement is used (Appendix B). The photons are produced isotropically from a vertical track at their assigned xyx-y locations, spanning between the liquid surface and the nominal anode. Geant4 optical physics is used to track the photons until they are absorbed by the photosoensors or other non-sensitive surfaces. For convenience we assume a sensor quantum efficiency (QE) of 1, but we will present the results as a function of detected photon number, so a <<1 QE only modestly affects the results by introducing an additional source of binomial fluctuations in the event-by-event photon count in a given sensor.
@@
Because of the symmetry provided by the circular shape of the photosensors and their hexagonal packing, the probability for a photon to be detected by a sensor centered at a radial distance of $r$ relative to the photon source in the horizontal plane is mostly determined by the value of $r$. Figure. 2 (left) shows the number of photons detected by 2" sensors, as a function of $r$, when 10,000 photons are emitted from the source. The probability decreases monotonically due to the inverse square law for photon flux, the decreasing solid angle subtended by the sensor, and the angular acceptance at the sensor surface. For small $r$ values, the diameter of the sensor dominates the aforementioned effects, with a relatively mild $r$ dependence within a sensor radius, beyond which a sharp drop in probability is observed. The different data sets correspond to different angular acceptance assumptions, which primarily impact photons hitting faraway sensors at large incident angles. For the following work, we choose a maximum photon acceptance angle of 85 degrees, motivated by the $\sim$50% Fresnel reflectance of quartz at this angle.
@@
Figure 2: Left - the simulated number of photons detected by 2" PMTs, as a function of the radial distance between the PMT center and the light source, for every 10,000 emitted photons; different colors represent different maximum photon acceptance angles (black-90 degree, red-85 degree, blue-80 degree); Right - mean photon detection probability as a function of radial distance, along with their best fit functions, for individual sensors of different sizes (from top to bottom: 3", 2", 1", 15 mm and 6 mm).
@@
Hereafter, we approximate the simulated photon detection probability as a function of the lateral distance $r$, referred to as the light response function (LRF), with an analytical form:
@@ 2.2 Fisher Information and Cramer-Rao Lower Bound
where the expectation is calculated over the observable space $X$. Under certain regularity conditions, the FI matrix may also be equivalently constructed as
@@
When there is a single parameter to be estimated, CRLB simply means that the smallest variance on the obtained parameter value is the inverse of the second derivative of the log-likelihood. Therefore, $\sqrt{[I^{-1}(\Theta)]_{ii}}$, or the Cramer-Rao Lower Bound (CRLB), is a limit of the precision that could be achieved with an unbiased estimator for $\hat{\theta_{i}}$ given the system response. In practice, maximizing the likelihood value, or minimizing the chi-square value or other metrics, for a given data set can provide an estimate of the parameters of interest. However, different estimators carry different uncertainties on the parameter values obtained, which may also depend on the specific implementation and initialization of the estimator. The smallest uncertainty in $\hat{\theta}$ from any unbiased estimator can be estimated by constructing and inverting the FI matrix.
@@
where $\eta_{k}(\Theta)$ is the Poissonian mean in the $k$th photosensor. As discussed in Sec. 2.1, we assume that the Poissonian mean in photosensor $k$ is completely determined by the number of photons emitted and the lateral distance between the light source and the sensor center, with a formula following Eq. 2.1 and parameters determined by optical simulations (Fig. 2). In our model, we ignore dark counts and other instrumental noise from the photosensors, which is justified by the capability of PMTs and SiPMs to definitively detect single photons and by their low dark count rates at cryogenic temperatures.
For the case of SS events in our TPC model, the parameters of interest are $\Theta=\{m_{1},x_{1},y_{1}\}$, corresponding to the number of photons emitted and the $x$ and $y$ coordinates of the event, and the observables are the number of photons detected by each photosensor $X=\{n_{k}\}$. Equations 2.1 and 2.6 can then be used to compute the FI matrix, and the CRLB values on these parameters can be derived accordingly. For the study of double-site (MS2) events, we add three similar parameters to describe the number of photons ($m_{2}$) and xyx-y location for the second vertex. However, to conveniently evaluate the detector’s sensitivity to the distance between the two vertices, instead of using the xyx-y coordinates of the second vertex, we use its distance from the first vertex ($d$) and the relative azimuth angle ($\phi$). Therefore, the additional parameters become $\{m_{2},d,\phi\}$ so that the CRLB on $d$ can be directly evaluated. In addition, for small $d$ values, a strong degeneracy between the photon counts at the two vertices, $m_{1}$ and $m_{2}$, is present; i.e., a decrease in $m_{1}$ can be approximately compensated for by increasing $m_{2}$ by a similar amount. To mitigate this issue, the final parameter set of $\Theta=\{m,x_{1},y_{1},f_{2},d,\phi\}$ is used, with $m=m_{1}+m_{2}$ and $f_{2}=m_{2}/m$.
This theoretical work is motivated by the studies of the “resolution measure” in optical microscopy, which investigates the smallest resolvable separation between two lit objects [19]; a similar study for microscopy can be found in [20]. For interested readers, we provide a detailed derivation of the FI matrix for independent Poisson distributions with binned data in Sec. A of the appendix.
2.3 TPC resolution predictions

### v3

@@ Abstract
Dual-phase xenon time projection chambers (TPCs) are widely used in searches for rare dark matter and neutrino interactions, in part because of their excellent position reconstruction capability in 3D. Despite their millimeter-scale resolution along the charge drift axis, xenon TPCs face challenges in resolving single-site (SS) and multi-site (MS) interactions in the transverse plane. In this paper, we build a generic TPC model with an idealized signal readout, and use Fisher Information (FI) to predict its theoretical capability of differentiating SS and MS events using the electroluminescence signal. We also demonstrate via simulation that, when only statistical photon noise is present, the theoretical limits can be approached with conventional reconstruction algorithms like maximum likelihood estimation, and with a convolutional neural network classifier. The implications of this study on future TPC experiments will be discussed.
@@ 1 Introduction
The time projection chamber (TPC) [1] is one of the most sensitive particle detection technologies. A TPC is a charge drift chamber that transports ionization electrons produced by particle interaction across the detector volume with an electric field; at the end of the detector the electrons are collected with a position sensitive charge sensing device. The charge sensor provides two-dimensional (2D) position information in the collection plane perpendicular to the electric field. In addition, if the interaction time is known, for example, via an accelerator beam signal or prompt scintillation accompanying ionization production, the third coordinate can be reconstructed from the electron drift time, enabling full three-dimensional (3D) event reconstruction. For gas TPCs that produce relatively long particle tracks, the track ionization density can be measured, allowing the particle type to be determined. For scintillating TPCs, the charge-to-scillation ratio may also help identify different particle interactions [2]. In addition, combining the scintillation and ionization energy channels enables an accurate energy reconstruction for the interaction by reducing the inter-channel fluctuation [3].
TPCs using liquid argon and xenon, which combine the benefits of a bright scintillator, a large target mass, and low electron affinity, are widely used in the search for rare particle interactions such as those expected from neutrinos and dark matter [4, 5, 6, 7, 8, 9]. A dual-phase noble liquid TPC further enhances the TPC’s sensitivity to low-energy interactions by introducing a thin gas layer above a liquid target. When ionization electrons are drifted into the gas, they can produce a large number of electroluminescence photons under a strong electric field. Therefore, both scintillation and ionization signals can be conveniently collected with a set of photosensors. Typically, an extracted electron can produce hundreds of electroluminescence photons, allowing signals as small as single ionization electrons to be detected with high efficiency [10]. Dual-phase argon and xenon TPCs have been widely used in dark matter search experiments that primarily focus on very low-energy signals. Currently, there are three xenon TPCs with multi-ton target masses in operation, including the LZ experiment [4] at the Sanford Underground Research Facility (SURF), the XENONnT experiment [11] at Laboratori Nazionali del Gran Sasso (LNGS), Italy, and the PandaX-nT experiment [8] at the Jinping Underground Laboratory in China. These detectors have been leading the searches for a range of hypothetical dark matter interactions, and their successes have inspired even larger detectors such as XLZD [12] and PandaX-20T to be proposed.
@@
One important application of SS/MS discrimination is in the search for neutrino-less double beta decays ($0\nu\beta\beta$) of the isotope 136Xe in a dual-phase xenon TPC [14]. $0\nu\beta\beta$ is predicted to occur in some double beta decay isotopes if neutrinos are Majorana fermions. A 136Xe $0\nu\beta\beta$ signal consists of two $\beta$ electrons whose energies sum to exactly $Q_{\beta\beta}=2.457$ MeV, while the primary background is induced by $\gamma$ rays from trace amounts of natural radioactivity in the detector materials. In liquid xenon, MeV-electrons are stopped within millimeters and the events usually appear point-like, while gamma rays tend to Compton scatter and deposit energy in multiple locations separated by several centimeters. Reconstruction of event topology therefore provides a powerful means of background discrimination. The nEXO experiment proposes to search for $0\nu\beta\beta$ in 136Xe with $\sim$5 tons of enriched xenon in a single-phase TPC and is expected to fully cover the inverted hierarchy region of neutrino mass with 10 years of operation [14]. Because of its large total mass (60–80 ton), the dual-phase TPC of XLZD would contain a comparable mass of 136Xe to that of nEXO and its sub-percent energy resolution near the 136Xe $Q$-value enables it to carry out a sensitive $0\nu\beta\beta$ search [15]. In both cases, 214Bi decay gamma rays are expected to be the dominant background in $0\nu\beta\beta$ energy region even with meticulous material cleanliness protocols [14, 15]. Therefore, a thorough understanding of the dual-phase xenon TPC’s capability in separating SS and MS events is critical.
@@
2 Detector model and resolution predictions
We develop a generic dual-phase TPC model with a light-based readout of ionization signals to study its theoretical performance in reconstructing SS and MS events across different event topologies as a function of the photosensor size and the total detected photon count. Photosensor diameters of 3", 2", 1", 15 mm, and 6 mm are considered, which represent the sizes of commonly used PMTs and SiPMs in dual-phase TPCs. Poissonian photon fluctuations are accounted for in this theoretical study but instrumentation effects that may introduce nonlinear signal response or additional noise are not considered. The detector model assumes direct photon detection with photonsensors, so the obtained results do not apply to liquid argon TPCs that use surface wavelength shifter coatings such as DarkSide50 [18], although an argon TPC model can be constructed and studied using a similar method.
@@ 2.1 Detector model
Figure 1: Left - an illustration of the generic TPC model used in our studies; Right - a cross-section view of a 40 cm TPC illustrating the photosensor array configuration; the asterisks represent two light emission sites separated by a distance $d$ and $r_{1}$ and $r_{2}$ represent the radial distances between the light sources and the PMT under investigation.
@@
The detector model, as illustrated in Fig. 1, is simplified to capture the essential features of a dual-phase xenon TPC. The electroluminescence region is modeled as a 1 cm xenon gas layer above a liquid–gas interface, where the electric field is assumed to be perfectly perpendicular to the liquid surface. With this approximation, electroluminescence photons excited by extracted electrons at a certain xyx-y location are produced uniformly along the vertical drift paths in this gap. Our model omits physical TPC electrodes, which generate the TPC drift field but may introduce field non-uniformities near the grid surfaces and cause shadowing effect in the light distribution. An array of photonsensors is placed 2 cm above the liquid surface to create a 1 cm separation between the upper end of photon tracks and the nearest sensor. This light source-sensor separation is smaller than those used in large TPCs, but is chosen to allow the studies of small sensors with the same TPC configuration; the effects of the sensor separation are discussed in Secs. 3.1.1 and 4.2. Each sensor is assumed to be circular in shape  [^1] SiPMs are often fabricated in square shapes, which can be modeled using the same method as described in this work; we model all sensors in the same shape to directly compare their performances. in a hexagonal close-packed configuration, yielding a photocathode fill factor of $\sim$91%. The sensor array is assumed to be sufficiently large to capture essentially all photons that could be collected, simulating the scenario of near-center events in a large TPC. While partial photon collection at detector edges could be implemented for a specific TPC geometry, such effects are not considered in this theoretical study. In practice, significant Fresnel reflections at large incident angles, combined with the small vertical separation between light source and sensor, limits the acceptance of photons from large radial distances. When photons are emitted from the center of the TPC, a 40 cm sensor array is able to capture virtually all detectable photons. Studies requiring light production far away from the center use larger detector sizes accordingly.
@@
A Geant4-based Monte Carlo simulation of the optical process is used to predict the photon detection pattern in the photosensor array for electroluminescence light emitted from a specific xyx-y location. In the simulation, all material surfaces, except the liquid-gas boundary, are assumed to be photon-absorbing so that no reflectance of light is in effect except at the liquid surface. Despite the simplification, our simulation approximately reproduces the radial photon distribution in the LUX experiment [13] when the same sensor placement is used (Sec. 3.1.1). The photons are produced isotropically from a vertical track at their assigned xyx-y locations, spanning between the liquid surface and the nominal anode. Geant4 optical physics is used to track the photons until they are absorbed by the photosoensors or other non-sensitive surfaces. For convenience we assume a sensor quantum efficiency (QE) of 1, but we will report all results as a function of detected photon number, so a <<1 QE only modestly affects the results by introducing an additional fluctuation in the event-by-event photon count in a given sensor. Similarly, saturation effects and instrumentation noise are not included in this study.
@@
Because of the symmetry provided by the circular shape of the photosensors and their hexagonal packing, the probability for a photon to be detected by a sensor is mostly determined by the horizontal distance between the photon source and the sensor center ($r$). Figure. 2 (left) shows the number of photons detected by 2" sensors, as a function of $r$, when 10,000 photons are emitted from the source. The probability decreases monotonically due to the inverse square law for photon flux, the decreasing solid angle subtended by the sensor, and the angular acceptance at the sensor surface. For small $r$ values, the diameter of the sensor dominates the aforementioned effects, with a relatively mild $r$ dependence within a sensor radius, beyond which a sharp drop in probability is observed. The different data sets correspond to different angular acceptance assumptions, which primarily impact photons hitting faraway sensors at large incident angles. For the following work, we choose a maximum photon acceptance angle of 85 degrees, motivated by the $\sim$50% Fresnel reflectance of quartz at this angle.
@@
Figure 2: Left - the simulated number of photons detected by 2" PMTs, as a function of the radial distance between the PMT center and the light source, for every 10,000 emitted photons; different colors represent different maximum photon acceptance angles (black-90 degree, red-85 degree, blue-80 degree); Right - mean photon detection probability as a function of radial distance (85 degree cutoff), along with their best fit functions, for individual sensors of different sizes (from top to bottom: 3", 2", 1", 15 mm and 6 mm).
@@
Hereafter, we approximate the simulated photon detection probability as a function of the radial distance $r$, referred to as the light response function (LRF), with an analytical form:
@@ 2.2 Fisher Information and Cramer-Rao Lower Bound
where the expectation is calculated over the observable space $X$. Under some regularity conditions, the FI matrix may also be equivalently constructed as
@@
When there is a single parameter to be estimated, Eq. 2.5 simply means that the smallest variance on the obtained parameter value is the inverse of the second derivative of the log-likelihood. Therefore, $\sqrt{[I^{-1}(\Theta)]_{ii}}$, or the Cramer-Rao Lower Bound (CRLB), is a limit of the precision that could be achieved with an unbiased estimator for $\hat{\theta_{i}}$ given the system response. In practice, maximizing the likelihood value, or minimizing the chi-square value or other metrics, for a given data set can provide an estimate of the parameters of interest. However, different estimators carry different uncertainties on the parameter values obtained, which may also depend on the specific implementation and initialization of the estimator. The smallest uncertainty in $\hat{\theta}$ from any unbiased estimator can be estimated by constructing and inverting the FI matrix.
@@
where $\eta_{k}(\Theta)$ is the Poissonian mean in the $k$th photosensor. As discussed in Sec. 2.1, we assume that the Poissonian mean in photosensor $k$ is completely determined by the number of photons emitted and the lateral distance between the light source and the sensor center, with a formula following Eq. 2.1 and parameters determined by optical simulations (Fig. 2). In our model, we ignore dark counts and read noise from the photosensors, which is justified by the capability of PMTs and SiPMs to definitively detect single photons and by their low dark count rates at cryogenic temperatures.
For the case of SS events in our TPC model, the parameters of interest are $\Theta=\{m_{1},x_{1},y_{1}\}$, corresponding to the number of photons emitted and the $x$ and $y$ coordinates of the event, and the observables are the number of photons detected by each photosensor $X=\{n_{k}\}$. Equations 2.1 and 2.6 can then be used to compute the FI matrix, and the CRLB values on these parameters can be derived accordingly. For the study of double-site events (MS with 2 vertices, dubbed MS2), we add three similar parameters to describe the number of photons ($m_{2}$) and xyx-y location for the second vertex. However, to conveniently evaluate the detector’s sensitivity to the distance between the two vertices, instead of using the xyx-y coordinates of the second vertex, we use its distance from the first vertex ($d$) and the relative azimuth angle ($\phi$). Therefore, the additional parameters become $\{m_{2},d,\phi\}$ so that the CRLB on $d$ can be directly evaluated. In addition, for small $d$ values, a strong degeneracy between the photon counts at the two vertices, $m_{1}$ and $m_{2}$, is present; i.e., a decrease in $m_{1}$ can be approximately compensated for by increasing $m_{2}$ by a similar amount. To mitigate this issue, the final parameter set of $\Theta=\{m,x_{1},y_{1},f_{2},d,\phi\}$ is used, with $m=m_{1}+m_{2}$ and $f_{2}=m_{2}/m$.
This theoretical work is motivated by the studies of the “resolution measure” in optical microscopy, which investigates the smallest resolvable separation between two lit objects [20]; a similar study for microscopy can be found in [21]. For interested readers, we provide a detailed derivation of the FI matrix for independent Poisson distributions with binned data in Appendix A.
2.3 Position resolution predictions